### v2

# ASTRID - An Automated and Scalable TRIaD for the Evaluation of RAG-based Clinical Question Answering Systems

Abstract
Large Language Models (LLMs) have shown impressive potential in clinical question answering (QA), with Retrieval Augmented Generation (RAG) emerging as a leading approach for ensuring the factual accuracy of model responses. However, current automated RAG metrics perform poorly in clinical and conversational use cases. Using clinical human evaluations of responses is expensive, unscalable, and not conducive to the continuous iterative development of RAG systems. To address these challenges, we introduce ASTRID - an Automated and Scalable TRIaD for evaluating clinical QA systems leveraging RAG - consisting of three metrics: Context Relevance (CR), Refusal Accuracy (RA), and Conversational Faithfulness (CF). Our novel evaluation metric, CF, is designed to better capture the faithfulness of a model’s response to the knowledge base without penalizing conversational elements. Additionally, our metric RA captures the refusal to address questions outside of the system’s scope of practice. To validate our triad, we curate a dataset of over 200 real-world patient questions posed to an LLM-based QA agent during surgical follow-up for cataract surgery - the highest volume operation in the world - augmented with clinician-selected questions for emergency, and clinical and non-clinical out-of-domain scenarios. We demonstrate that CF predicts human ratings of faithfulness more accurately than existing definitions in conversational settings. Furthermore, using eight different LLMs, we demonstrate that the three metrics can closely agree with human evaluations, highlighting the potential of these metrics for use in LLM-driven automated evaluation pipelines. Finally, we show that evaluation using our triad of CF, RA, and CR exhibits alignment with clinician assessment for inappropriate, harmful, or unhelpful responses. We also publish the prompts and datasets for these experiments, providing valuable resources for further research and development.
ASTRID - An Automated and Scalable TRIaD for the Evaluation of RAG-based Clinical Question Answering Systems
Yajie Vera He [^2] Authors contributed equally. Ufonia Limited yh@ufonia.com                        Mohita Chowdhury [^2] Authors contributed equally. [^1] Now at Google DeepMind. Ufonia Limited chowdhury.mohita@gmail.com                        Jared Joselowitz Ufonia Limited jj@ufonia.com
Aisling Higham Ufonia Limited ah@ufonia.com                        Ernest Lim Ufonia Limited / University of York el@ufonia.com
1 Introduction
The healthcare industry is increasingly adopting automation to meet rising demands on resources Ruiz and Duffy (2021). LLMs, due to their capabilities, have become increasingly popular in supportive clinical applications such as note-taking and summarization Cascella et al. (2023). A crucial aspect of patient care is the ability to ask questions and receive answers, which has been enhanced by advancements in QA systems powered by LLMs. However, the issue of hallucination remains a significant barrier to using LLMs for clinical QA systems Rawte et al. (2023). RAG is a technique developed to address hallucination and ensure context appropriateness Lewis et al. (2020). Despite these advancements, RAG systems lack sufficient evaluation metrics and frameworks, making it difficult to quantitatively establish their safety and identify system deficiencies. Figure 1 illustrates the limitations of current clinical evaluation approaches and how automated methods address these challenges. This work examines evaluation limitations and applies safety engineering to identify hazard cases in clinical QA Hawkins et al. (2022); Ericson et al. (2015). We develop a robust, scalable framework of metrics to systematically demonstrate how developers can mitigate potential hazards in LLM-based QA systems for clinical use. Using real patient questions from clinical trials on cataract post-operative recovery, we illustrate how these metrics can be interpreted in a clinical context. We validate our metrics by proving they model human ratings better than previous metrics and effectively predict clinical harm, usefulness, and inappropriateness as labeled by specialist doctors. Our aim is to establish a foundation for developing and assessing LLM-powered clinical QA systems and encourage further research in this area.
Figure 1: Clinical human evaluation is the gold standard for clinical relevance, but is inherently unscalable. Current automated RAG evaluation metrics are not suited for clinical or conversational contexts. We propose ASTRID to address these limitations towards scalable, and clinically relevant evaluation of RAG-based Clinical QA systems.
Our contributions are summarized as follows:
  • A hazard analysis of clinical QA systems inspired by safety engineering principles.
  • A new suite of metrics for clinical QA systems motivated by this analysis.
  • A formal evaluation of these metrics and their alignment with human ratings.
  • An analysis of how these metrics collectively predict clinical harm, usefulness, and inappropriateness with high accuracy.
  • An assessment of how these metrics can be automated across eight different LLMs.
2 Related work
2.1 Background to clinical QA evaluation
Clinical QA systems powered by LLMs have generated significant recent interest. Already, some LLMs have demonstrated capabilities to generate more accurate responses Thirunavukarasu et al. (2023); Bernstein et al. (2023); Samaan et al. (2023); Xie et al. (2023); Van Bulck and Moons (2024), and sometimes even more empathetic than doctors across various clinical contexts Lee et al. (2024). However, LLMs can generate plausible-sounding, but factually incorrect responses, commonly referred to as "hallucinations" Ji et al. (2023). Moreover, LLMs have knowledge cut-off date Ovadia et al. (2023) and this can pose significant safety risks in healthcare. While these issues can be mitigated using RAG, risks still remain.
To address some of these risks specific to clinical QA systems using RAG, various efforts have been made to develop performance benchmarks. Currently, published benchmarks often utilize multiple-choice or categorical ground-truth answers for responses Xiong et al. (2024); Li et al. (2024); Wu et al. (2024); Nori et al. (2023), which fail to capture the complexities and risks associated with open-ended response generations. Where open-ended answers are evaluated, n-gram-based metrics such as BLEU Papineni et al. (2002) or ROUGE Lin (2004), historically used for machine translation, have been used Chen et al. (2019). However, these evaluations have been criticized for failing to capture the nuanced requirements of clinical QA, and even transformer-based metrics such as BertScore Zhang et al. (2019) have faced numerous semantic limitations Dada et al. (2024).
A key feature of these risks in the context of open-ended clinical QA is their non-binary nature (i.e., an answer is not simply "safe" or "unsafe" on a single axis). Consequently, the gold standard for assessing clinical inappropriateness remains human evaluation. For instance, Google’s work in clinical QA involved both clinicians and lay individuals, labeling responses based on various axes such as the likelihood and severity of harm, alignment with scientific consensus, and helpfulness Singhal et al. (2023a). Similarly, other studies have employed multi-axis evaluations with human clinicians to assess the overall appropriateness of responses for open-ended clinical QA Mukherjee et al. (2024); Singhal et al. (2023b); Zakka et al. (2024); Chowdhury et al. (2023).
However, this approach is highly unscalable due to the significant time and resources required for continuous human evaluation with specialist clinicians. Additionally, large end-to-end question-output evaluations hinder iterative development and rapid prototyping of RAG-based clinical QA systems, as they often fail to provide clear guidance to developers on how to adapt their RAG pipelines to resolve clinical performance issues.
2.2 Current RAG metrics
Evaluating RAG systems presents challenges due to their hybrid structure and the overall quality of the output often depends on multiple components within these systems. While attempts have been made to assess the overall quality of responses using deterministic methods Liu et al. (2023); Lyu et al. (2024), most current evaluation metrics for RAG systems use an ensemble of component-level assessments, the majority of which leverage LLMs as judges Yu et al. (2024). Broadly, RAG pipelines and the axes used to assess their performance can be broken down into the following components.
Retrieval component
The retrieval component is responsible for extracting relevant context from a knowledge source to match a given query.
  • Relevance (Context $\leftrightarrow$ Question): Measures how well the retrieved context matches the query’s information needs.
  • Accuracy (Relevant Context $\leftrightarrow$ Context Candidates): Assesses the accuracy of retrieved context compared to a set of candidates.
Generation component
where a model generates a response using information from the given context
  • Relevance (Response $\leftrightarrow$ Question): Evaluates the alignment of generated responses with the question’s intent.
  • Faithfulness / Groundedness (Response $\leftrightarrow$ Context): Assesses whether generated responses accurately reflect the retrieved context.
  • Correctness (Response $\leftrightarrow$ Sample Response): Measures the factual correctness of generated responses against a sample or standard response.
These component evaluations have been variably implemented with popular tools including TruEra’s RAG Triad Trulens (2023), and LangChain Bench LangChain . Additionally, LLM-as-a-judge-based frameworks like RAGAS Es et al. (2023), and ARES Saad-Falcon et al. (2023) have popularized common evaluation triads to capture possible permutations of the above components. Refer to Appendix A for an example of how the three components of the RAG system can be judged by LLMs, using RAGAS as an example.
2.2.1 Limitations of current metrics
Faithfulness
The established methods to measure Faithfulness break down a model’s response into granular statements and then evaluate each statement’s consistency with the context Es et al. (2023). This approach aims to create more focused assertions that consider the context of both the question and the answer. It is particularly advantageous when answers are short and lack sufficient context when reviewed in isolation, as demonstrated in Figure 7. However, in the context of clinical conversations, this approach has the following shortcomings:
  • Summarizing the response into statements often neglects the clinical nuances in the original response (Figure 8).
  • Creating statements from both the patient’s question and the agent’s answer prevents the independent review of the agent’s answer concerning the context. This is especially problematic when the combined statement contains factually incorrect information (Figure 9).
  • Dialogue agents, particularly in clinical settings, are prompted to respond empathetically and conversationally. Statements constructed from the agent’s acknowledgments and questions, such as those meant to clarify or follow up on the patient’s queries or concerns, are penalized by existing faithfulness definitions (Figure 10).
Answer Relevance
Evaluating answer relevance is critical in QA systems to ensure generated responses align with query’s intent. However, most current definitions focus on lexical or semantic similarity between the question and the response [Siriwardhana et al. (2023); Es et al. (2023)]. This has a number of drawbacks:
  • It overemphasizes surface-level topic matching without accounting for deeper contextual understanding.
  • It fails to account for whether a context is appropriate given a clinical context.
  • It does not handle "non-answers", meaning it struggles to determine when a system correctly discerns that a question falls outside its scope of relevance or when there is insufficient information to provide a safe and accurate response.
Furthermore, these metrics often do not reflect when the system appropriately refuses to address the question. This is critical in a clinical setting as clinicians, and similarly clinical QA systems must stay within their scope of practice.
3 Proposed approach
3.1 Deriving metrics towards a safety case
In order to align our framework towards the evidence required to demonstrate if a clinical system is safe, we sought inspiration from published safety engineering frameworks - namely the Safety Assurance of autonomous systems in Complex Environments (SACE) guidance Hawkins et al. (2022). Structured safety engineering approaches have been applied towards the assurance of high-integrity autonomous systems (AS) such as maritime vessels Nakashima et al. (2023), automotive Rahman et al. (2023); Hunter et al. (2024), aerospace Torens et al. (2022), and healthcare domains Jia et al. (2022); Festor et al. (2022).
The SACE framework, in particular, provides a process to systematically integrate safety assurance into the development of AS whilst considering the system and its environment. Whilst we do not report all artifacts from the process in its entirety, we highlight a few key steps in this process that have been applied towards ASTRID’s design. Specifically, we considered the principles of:
  • Operating Context Assurance: Identifying the different clinical scenarios a patient might pose to a QA agent (see Figure 2).
  • Hazardous Scenario Identification: Analyzing how RAG systems could behave in hazardous ways within these scenarios.
  • Safe Operating Concept Assurance: Defining how an ideal system should behave in response to different queries.
  • Out-of-Context Operation Assurance: Determining how the system should respond when a question falls outside the scope.
Figure 2: Clinical Operating Contexts that face a clinical QA agent.
The clinical context is essential for determining appropriateness. For instance, the question, "Is it normal to have stomach cramps and vomiting?" would be irrelevant during a follow-up appointment for routine eye surgery, and the system should not to respond. However, if posed by a patient who has just been discharged after bowel surgery, it is not only highly relevant but also critical that the system provides a response. This is illustrated in Figure 3.
These concepts were outlined in a workshop where the dataset of real-world questions posed by patients to a voice-based AI conversational question were reviewed. The workshop consisted of AI developers, a clinician, and a safety practitioner (summarized in Figure 14). The analysis served as a bridge between subjective clinical assessments of harm and helpfulness and component-level validation scenarios for appropriate system performance. From the subsequent hazard analysis and the definition of safe operating requirements, it became clear that existing RAG evaluation metrics do not adequately capture key clinical risks in a conversational QA setting.
Figure 3: Whether questions are clinically appropriate relies heavily on the clinical context, thus metrics need to be situated in this context.
3.2 A novel set of metrics and a framework to assess safety risks
Current RAG QA metrics do not correlate to clinical risks, and have varying levels of validation against human evaluations, often performing poorly in conversational contexts. To our knowledge, no efforts have been made to connect QA system performance—measured by these metrics—with real-world clinician grading of clinical harms. For developers to meaningfully assess whether a clinical RAG QA system meets safe operating concepts, a framework is needed that is validated for clinical use, scalable, and capable of accounting for nuanced clinical contexts.
We propose a novel Automated and Scalable TRIaD (ASTRID) analysis framework for RAG-based clinical QA systems. ASTRID consists of three reference-free LLM-based metrics: Refusal Accuracy (RA), Conversational Faithfulness (CF) and Context Relevance (CR) (Figure 4). In the subsequent sections, we will illustrate how we validate each of the metrics and the overall framework based on a real world data from patients speaking to clinical conversational agents, augmented to ensure sufficient test case coverage.
Figure 4: ASTRID - an Automated and Scalable TRIaD for evaluating clinical QA systems leveraging RAG - consisting of three metrics: Context Relevance, Refusal Accuracy, and Conversational Faithfulness assessed within a clinical context.
3.2.1 Conversational Faithfulness (CF)
Evaluating the alignment of a response with the provided information is crucial for QA systems using RAG. Existing metrics fail to capture the complexities of conversational agents in clinical settings. We propose a newly-defined metric, Conversational Faithfulness, to address this gap.
Given an answer-context pair, CF is defined as the proportion of information-containing sentences that are faithful to the context. To calculate CF, we employ the following steps:
  1. 1.
    We categorize sentences in the response as either "information-containing" or "not information-containing" and extract the information-containing sentences.
  2. 2.
    We determine whether the information-containing sentences are grounded in context.
The prompts used to execute these steps are provided in Appendix 11. Finally, CF is calculated as follows:
$\texttt{CF}=\begin{cases}1,&\text{if }N=0\\ \frac{Y}{N},&\text{if }R\geq 0\\ 0,&\text{otherwise}\end{cases}$ (1)
where:
Y = Number of information-containing statements grounded in context
N = Total number of information-containing statements
R = N - (Y + Number of information-containing statements not grounded in context")
3.2.2 Refusal Accuracy (RA)
As discussed in previous sections, ensuring that a QA system appropriately declines to respond when a question is unanswerable or contextually inappropriate is critical in clinical settings, particularly for LLM-powered systems prone to generating ungrounded responses. Existing metrics do not capture this behavior, prompting us to introduce Refusal Accuracy.
Refusal Accuracy measures a system’s ability to withhold a response when no relevant information is available. It is evaluated using binary labels indicating whether the system appropriately refuses to answer. The prompt used for this assessment is provided in Appendix 13.
3.2.3 Context Relevance (CR)
For clinical QA systems built on RAG, using the right context when generating responses is essential. This is typically achieved by creating embeddings of the query and the knowledge source, which are then passed through a retriever Lewis et al. (2020); Ding et al. (2024). The retrieval component of a RAG system takes in the encoded query and retrieves the top matches from the knowledge source index, which is then passed to the LLM agent as context Salemi and Zamani (2024).
For voice-based conversational QA systems, our dataset analysis indicates that user queries typically consist of no more than two questions per turn. Additionally, specialised knowledge sources within a specific clinical scope are relatively small and focused, in contrast to the extensive databases used for general clinical QA. Given that multiple pieces of information may be necessary to comprehensively answer a query in this dataset, the clinical RAG QA system employed in this evaluation retrieves the top three context chunks.
Many existing CR definitions penalise additional retrieved contexts Es et al. (2023); Saad-Falcon et al. (2023). However, we place greater emphasis on measuring the completeness of clinical information present in the retrieved context. To better suit our use case, we simplify the CR definition and define it as a binary label indicating whether the retrieved context is relevant to the query. Appendix 12 shows the prompt used to achieved this.
4 Method
We conduct several experiments using datasets sourced from real clinicians and open-source datasets to support the following claims:
  1. 1.
    Our Conversational Faithfulness metric models human Perceived Faithfulness (PF) more accurately than existing definitions.
  2. 2.
    ASTRID can predict clinician ratings of harmfulness, helpfulness, and inappropriateness.
  3. 3.
    ASTRID is straightforward for LLMs to use, making them highly automatable.
4.1 Dataset Curation
We curated datasets to support evaluation of our proposed metrics using both real-world patient interactions and augmented clinical questions.
Real-world Data Collection.
We collected real-world patient questions using an autonomous telemedicine assistant capable of conducting phone conversations and responding to recovery-related queries following cataract surgery. From interactions with 120 patients at two UK hospitals, we extracted 102 unique patient questions. All patients provided explicit verbal consent for anonymized data usage. This setup enabled collection of naturally occurring, often noisy inputs (e.g., mistranscriptions, statements, or truncated queries), representative of real-world deployments.
Response Generation.
To generate responses to these queries, we curated a domain-specific knowledge base on cataract surgery with two ophthalmic surgeons. We then employed three LLMs—PaLM-2 (text-bison@002, Anil et al. (2023)), Mistral-7B Jiang et al. (2023), and LLaMA-8B Touvron et al. (2023)—within a RAG-based QA setup to produce answers. This yielded 306 question-answer-context triplets. We then filtered examples to retain only those exhibiting conversational elements (e.g., acknowledgments, follow-ups), resulting in a refined dataset of 206 triplets.
4.1.1 Balancing by Perceived Faithfulness
Two independent annotators assessed the Perceived Faithfulness of answers relative to their context, without being shown the original question. Disagreements were discussed and resolved by consensus. To ensure balanced training data, we sampled an equal number of perceived faithful and unfaithful responses, resulting in 148 triplets (74 faithful, 74 unfaithful).
4.1.2 Augmenting with Out-of-Scope Data
To evaluate the Refusal Accuracy metric, we added 45 out-of-scope questions selected by clinicians from the open-source HealthSearchQA dataset Singhal et al. (2023a). These were paired with answers generated using PaLM-2 and LLaMA-8B, resulting in 90 new triplets. Combined with our in-scope dataset, this yielded 238 triplets.
4.1.3 Human Labeling of Metrics
Independent annotators labeled the 238 triplets for both the older "statement-level" definition of Faithfulness and our proposed Conversational Faithfulness. Each (answer, context) pair was annotated using both definitions, and consensus was reached through discussion. The final dataset (we name FaithfulnessQAC) includes human ratings for faithfulness, CF, and PF.
4.1.4 Constructing UniqueQAC
To support Claim 2, we created a version of the dataset with unique questions by selecting 87 in-scope triplets from FaithfulnessQAC where each question was distinct. For the out-of-scope examples, we randomly sampled 45 triplets. Each triplet includes a question, its answer (randomly selected from the LLM outputs), and corresponding context. This resulted in the 132-example UniqueQAC dataset, labeled for CF, CR, and RA.
4.1.5 Constructing ClinicalQAC
To support Claim 3, we further enriched UniqueQAC by asking two ophthalmic surgeons to label each triplet across three axes:
  1. 1.
    Clinical Harm: Is the response potentially harmful?
  2. 2.
    Helpfulness: Does the response assist the patient?
  3. 3.
    Appropriateness: Is any content clinically inappropriate or incorrect?
Most original responses were safe and helpful, so we replaced a subset with responses authored by clinicians that intentionally included unhelpful, harmful, or inappropriate content to balance the dataset. The resulting 132-example dataset, ClinicalQAC, supports clinical risk evaluation. Figure 16 shows the final class distribution.
We provide further details of the dataset curation process in Appendix E.
4.2 Experiments
We break down this section by Claims 1, 2, and 3, detailing the different experiments we conducted to support them and discussing the results.
4.2.1 Demonstrating alignment of Conversational Faithfulness with human perception
Setup.
To demonstrate that our metric, Conversational Faithfulness, aligns more closely with human perception of faithfulness than previous definitions, we perform the following:
  1. 1.
    We treat CF as a diagnostic test that predicts human PF. We compare it with the classification based on the previous definition of faithfulness, which we call RF (inspired by RAGAS), and conduct a ROC analysis for both. To do this, we use human ratings of CF and PF from the FaithfulnessQAC dataset.
  2. 2.
    We use Pearson, Spearman, and Kendall Tau correlation coefficients to correlate human ratings of CF and RF with PF.
We used human-annotated CF and RF scores rather than LLM-generated ones, eliminating potential model artifacts from the analysis.
Results.
From Figure 5, we observe that our metric CF achieves an AUC of 0.98, outperforming RF (0.83), demonstrating superior alignment with human ratings.
The correlation analysis in Table 1 reinforces the advantages of CF. All three metrics reveal that CF exhibits significantly stronger alignment with human judgments ($\geq$ 0.84) compared to RF ($\leq$0.57). The consistency across these measures suggests that CF better captures both linear and rank-order relationships with human perceptions of faithfulness. These results demonstrate that CF outperforms existing faithfulness metrics in conversational contexts, offering a more reliable automated measure aligned with human judgment.
Figure 5: ROC curve for CF and RF. The ROC curve for CF has an area of 0.98 and the ROC curve for RF has an area of 0.83.
Correlation Type CF vs PF RF vs PF
Pearson correlation 0.90 0.57
Spearman correlation 0.90 0.57
Kendall Tau correlation 0.84 0.50
Table 1: Correlation coefficients for CF and RF against PF
Harmfulness Helpfulness Inappropriateness
Harmful Unharmful Helpful Unhelpful Yes Slightly No
RandomForest 0.82 0.80 0.73 0.70 0.67 0.00 0.78
SVM 0.86 0.86 0.73 0.70 0.67 0.00 0.78
Gaussian Naive Bayes 0.86 0.86 0.73 0.70 0.80 0.31 0.57
Neural Network 0.82 0.80 0.73 0.70 0.67 0.00 0.78
Average 0.84 0.83 0.73 0.70 0.71 0.08 0.73
Table 2: F1-scores when CF, CR, RA and scope of practice are used as features to predict Harmfulness, Helpfulness and Inappropriateness using different models.
Model CF CR RA
Accuracy F1 Accuracy F1 Accuracy F1
mistral-7B 0.44 ± 0.009 0.19 ± 0.021 0.73 ± 0.006 0.67 ± 0.010 0.91 ± 0.000 0.88 ± 0.000
llama-3-8B 0.39 ± 0.005 0.05 ± 0.009 0.52 ± 0.006 0.64 ± 0.005 0.90 ± 0.000 0.88 ± 0.000
llama-3.3-70B 0.45 ± 0.006 0.23 ± 0.015 0.82 ± 0.012 0.82 ± 0.013 0.90 ± 0.005 0.87 ± 0.006
mistral-large-2402 0.50 ± 0.006 0.33 ± 0.014 0.78 ± 0.006 0.72 ± 0.010 0.84 ± 0.009 0.81 ± 0.008
claude-3.5-sonnet 0.72 ± 0.004 0.74 ± 0.004 0.81 ± 0.004 0.81 ± 0.005 0.91 ± 0.000 0.88 ± 0.000
gemini-2-flash 0.69 ± 0.016 0.77 ± 0.012 0.75 ± 0.014 0.78 ± 0.010 0.93 ± 0.005 0.90 ± 0.007
gpt-4o 0.69 ± 0.011 0.74 ± 0.011 0.86 ± 0.012 0.85 ± 0.013 0.92 ± 0.006 0.89 ± 0.006
gpt-o3-mini 0.68 ± 0.017 0.69 ± 0.017 0.87 ± 0.009 0.87 ± 0.009 0.95 ± 0.010 0.94 ± 0.014
Table 3: LLM Performance on ASTRID Metrics (Accuracy & F1-Scores).
4.2.2 Predicting clinical assessments using our triad of metrics
Setup.
We investigated whether our three metrics (CF, CR, and RA) could effectively predict clinician assessments of QA system responses using the ClinicalQAC dataset. Our goal was to demonstrate that these automated metrics could identify potentially harmful, unhelpful, or inappropriate responses.
We reserved 17.5% of the dataset as a balanced test set (see Figure 6). We then manually choose triplets to ensure balanced categories and randomly sample 79% of the remaining dataset for the train split and use the remaining 21% as the val split.
We then train four models (Random Forest, SVM, Gaussian Naive Bayes, and Neural Network) to demonstrate how our triad can independently predict harmfulness, helpfulness, and inappropriateness when the scope of practice (within scope/out-of-scope) is taken into account. We subsequently test the results on the test set and report precision, recall and F1-scores.
Figure 6: ClinicalQAC test split distribution across categories
Results.
In Table 2, our experiments demonstrate that the triad of metrics serves as a strong predictor across all clinical assessment categories. Using our triad and the scope of practice, we can predict clinician rating of harmfulness with an average F1-score of 0.835. We can also predict helpfulness with an average F1-score of 0.715.
For inappropriateness prediction, the models showed strong performance in identifying clearly appropriate and inappropriate responses, with F1-scores of 0.73 for "No" and 0.70 for "Yes" classifications. However, detecting "slightly" inappropriate content proved challenging, with an average F1-score of 0.08. This difficulty aligns with human assessment patterns, as evidenced by the lower inter-annotator agreement (65%) for inappropriateness ratings prior to resolution. We report other inter-annotator scores in the appendix in Table 4.
Figure 15 provides illustrative examples demonstrating how these metrics can identify potentially harmful failure modes at the individual question level, offering developers a framework for correlating system behavior with clinician-assessed harms.
4.2.3 Automatability of our triad of metrics
Setup.
In this experiment, we assess whether LLMs can be used to automatically compute the three ASTRID metrics without requiring human annotation. Table 3 presents the performance of eight LLMs across these metrics, comparing their scores against human-labeled ground truth. The inference details for each LLM, including deployment environment and token limits, are provided in the Appendix in Table 5.
Results.
The results indicate that several LLMs achieve reasonably close agreement with human ratings, supporting the feasibility of automated metric evaluation. Notably, GPT-4o, claude-3.5-sonnet, and Gemini-2-Flash exhibit high CF F1-scores (0.74, 0.74, and 0.77 respectively), suggesting strong alignment with human judgments of conversational faithfulness. Similarly, GPT-o3-mini achieves the highest CR Accuracy (0.87) and RA Accuracy (0.95), demonstrating superior retrieval relevance and appropriate refusal behavior.
However, smaller models like LLaMA-3-8B and Mistral-7B struggle, particularly in CF, with F1-scores of 0.05 and 0.19, respectively, indicating difficulty in distinguishing context-grounded versus ungrounded responses. This suggests that larger, more advanced models may be better suited for automated ASTRID metric computation due to their improved reasoning and contextual understanding.
These findings suggest that LLMs can serve as scalable evaluators of RAG-based clinical QA systems, reducing reliance on human annotators. While further prompt engineering and fine-tuning may improve alignment with human ratings, the ability of models to automatically compute CF, CR, and RA offers a promising direction for developing continuous and automated evaluation pipelines for clinical LLM systems. Further breakdowns of model performance across all ASTRID metrics are presented in Appendix H.
5 Conclusion
We present ASTRID, an Automated and Scalable TRIaD for evaluating clinical QA systems leveraging RAG. ASTRID comprises three metrics — Conversational Faithfulness, Refusal Accuracy, and Context Relevance — designed to address the limitations of existing evaluation frameworks in clinical settings. Our experiments demonstrate that CF aligns more closely with human judgments of faithfulness compared to previous definitions, and our triad of metrics is the first to correlate system performance measures with clinician assessments of harmfulness, helpfulness, and inappropriateness with high accuracy. We also highlight the potential for these metrics to be automatable using current LLMs, making them suitable for iterative development and the continuous evaluation of clinical QA systems. By publishing our datasets and prompts, we aim to provide valuable resources for further research and development in the field. Future work should expand on end-to-end conversational evaluations and incorporate usability metrics to ensure a comprehensive assessment of clinical QA systems.
Limitations
One limitation of our approach is that we focus on single-turn safety rather than end-to-end conversations. End-to-end conversations introduce an additional element of decision-making and context-continuity that need to be assessed for a holistic evaluation of a QA system. Further work should explore multi-turn interactions to ensure comprehensive safety, reliability, and extended dialogue evaluation.
Our metrics and evaluation frameworks are centered around safety. Notably, we have not incorporated usability aspects into our evaluations, such as robustness to mistranscriptions Yu et al. (2024), measures of clinical empathy Sorin et al. (2023), latency, brevity, or user satisfactionMukherjee et al. (2024). Future research incorporating these dimensions will provide a more well-rounded assessment of QA systems in real-world clinical environments. Beyond single-turn interactions as well, future work should explore multi-turn interactions, and interactions adjacent to clinical question answering, such as information gathering. In parallel, we plan to extend the framework to cover multi-domain evaluations, adapting the framework to support a variety of clinical specialties across not just elective-type pathways.
A strength of the study was it utilized a real-world dataset of questions posed to a voice-based AI agent. This dataset included mistranscriptions, statements, and truncated questions, reflecting real-world scenarios. Additionally, we developed a clinician-generated dataset in the clinical domain of hip surgery follow-up to explore generalizability. However, we limited our analysis to the real-world dataset to focus on actual hazard cases rather than hypothetical ones. Due to the availability of data, we also only focused on one clinical use-case. Further research needs to expand these findings across other clinical settings beyond ophthalmology and with larger datasets. Nevertheless, we publish the augmented ClinicalQAC dataset, complete with responses and labels, for open use.
Ethics and Data Statement
The research presented in this article is fully consistent with the ACM Code of Ethics and Professional Conduct (https://www.acm.org/code-of-ethics). This paper does not involve crowd-sourcing or research with human subjects. We adhere to the UK Health Research Authority (HRA) guidelines on using health and care information for this work and completed the Medical Research Council (MRC) ethics review checklist, indicating that research ethics approval was not required.
The question data were obtained from an aggregated and anonymized pool collected during the routine deployment of a CE-marked clinical conversation Al agent in the UK between 2021-2022 (Dora, Ufonia Limited). All individuals gave explicit, documented verbal consent for anonymized data to be used for research purposes. Data use also complies with data protection guidelines in the UK. The study data is aggregated and anonymized with no markers for identifiability per the current Information Commissioner’s Office (ICO) code of practice, and HRA research glossary. Specifically, the dataset has none of the ‘key indicators of identifiability’, and we have ensured there is no risk of re-identification of any person from the data source. Audio data were not used for this study. This data is made available for quality and representativeness review purposes.
Demographic data are not available from this dataset, as the questions are disambiguated and not linked to individual patients due to the anonymization measures described above. Although representative of real routine actual use, we acknowledge that the lack of demographic data is a limitation. However, proxy measures of demographic details can be referenced from two published clinical trials conducted with a similar geographical and clinical patient cohort.
The HealthsearchQA dataset, which was used to augment our dataset, is published under a Creative Commons Attribution 4.0 International License (https://www.nature.com/articles/s41586-023-06291-2) and is freely available in its entirety.
GPT-4 was used only for spell-checking and grammatical proof reading of this manuscript via the OpenAI ChatGPT interface.
Appendix A RAGAS definitions: Context Relevance, Answer Relevance and Faithfulness
For example, the RAGAS triad has three components as judged by LLMs:
Context Relevance:
CR=number of relevant context sentencestotal number of $\texttt{CR}=\frac{\texttt{number of relevant context sentences}}{\texttt{totalnumber of sentences}}$ (2)
Answer Relevance:
Answer $\texttt{Answer Relevance}=\frac{1}{N}\sum_{i=1}^{N}\frac{E_{g_{i}}\cdot E_{o}}{\|E_{g_{i}}\|\|E_{o}\|}$ (3)
Where:
  • $E_{g_{i}}$ is the embedding of the generated question $i$.
  • $E_{o}$ is the embedding of the original question.
  • $N$ is the number of generated questions, which is 3 by default.
Faithfulness
$F=\frac{|V|}{|S|}$ (4)
Where:
  • $|V|$ is the number of statements that were supported according to the LLM.
  • $|S|$ is the total number of statements.
Appendix B Examples of limitation of current metrics
Figure 7: Example 1: Extracting statements is particularly advantageous when answers are short and lack context when reviewed in isolation
Figure 8: Example 2: Summarising the response into statements often neglects the clinical nuances in the original response.
Figure 9: Example 3: Creating statements from both the patient’s question and the agent’s answer prevents the independent review of the agent’s answer concerning the context. This is especially problematic when the combination contains factually incorrect information.
Figure 10: Example 4: Dialogue agents, particularly in clinical settings, are prompted to respond empathetically and conversationally.
Appendix C Prompts
Figure 11: Prompts for computing Conversational Faithfulness
Figure 12: Prompt for computing Context Relevance
Figure 13: Prompt for computing Refusal Accuracy
Appendix D Illustrative Examples of ASTRID
Figure 14: Hazards Analysis through applying SACE framework in the context of a clinical QA agent. Examples are int he context of a post-operative cataract surgery call.
Figure 15: Illustrative examples of ASTRID metrics and correlated clinician labels with both in-scope and out of scope questions. Potential approaches to improve on metrics are discussed in interpretation. Green boxes demonstrate expected metric outcomes for that context.
Appendix E Dataset Curation Process
To collect real-world patient questions, we used an autonomous telemedicine assistant capable of conducting phone conversations and answering patient questions regarding their recovery following cataract surgery. From these interactions, we gathered 102 unique questions from 120 patients from calls that took place as a standard of their care at two UK hospitals. All patients explicitly consented to the use of their anonymised data for research purposes.
To generate answers to these questions, we curated a knowledge source on cataract surgery with the help of two ophthalmic surgeons. We then employed three LLMs – Palm-2 (text-bison@002, Anil et al. (2023)), Mistral-7B [Jiang et al. (2023)) and Llama-8B [Touvron et al. (2023)] – as part of a RAG-based QA agent to generate responses to the 102 questions. This process resulted in a dataset of 306 question-answer-context triplets.
Subsequently, we sampled triplets where the answers included conversational elements such as acknowledgements and follow-up questions, reflecting real-world conversational responses. This refined dataset comprises 206 question-answer-context triplets.
E.1 Balancing by Perceived Faithfulness
We employed two independent labellers to assess "faithfulness" for the 206 examples by showing them only the answer and the context. We asked them to use their own judgment to determine whether a given answer was faithful to the context. We refer to this measure of human judgment as Perceived Faithfulness (PF). The labellers discussed and resolved any disagreements to ensure consensus.
To create a balanced dataset, we sample an equal number of perceived faithful and unfaithful responses. This process resulted in a dataset consisting of 74 faithful and 74 unfaithful responses, culminating in a total of 148 question-answer-context triplets.
E.2 Augmenting with out-of-scope data
For a holistic evaluation, we augmented this dataset with 45 out-of-scope questions selected by two clinicians from the open-source dataset HealthSearchQA [Singhal et al. (2023a)]. We created 90 question-answer-context triplets using the same process mentioned earlier with only Palm-2 and Llama-8B, resulting in a comprehensive dataset of total 238 triplets.
E.3 Human labeling
We employed two independent labellers to follow the older definition of Faithfulness and Conversational Faithfulness (CF) to generate labels for the (answer-context) pairs from the 238 examples. After the labellers resolved disagreements, we combined these ratings with the perceived faithfulness human ratings to create a comprehensive dataset of 238 question-answer-context triplets with human-rated faithfulness, CF, and PF. We name this dataset FaithfulnessQAC and make it available for research purposes.
E.4 Creating a dataset of unique questions
To support Claim (3), we sampled 87 out of 148 triplets where each question is unique, with the corresponding answers randomly selected from one of the three LLM responses. For the out-of-scope data, we randomly chose 45 triplets with responses from between the two models. The final dataset contains 132 question-answer-context triplets, and we name it UniqueQAC, making it available for research purposes.
E.5 Creating a dataset of unique questions
To support Claim (2), we need a dataset where each question is unique, as Context Relevance is measured between the question and the context. To achieve this, we sampled 87 out of 148 triplets where each question is unique, with the corresponding answers randomly selected from one of the three LLM responses. For the out-of-scope data, we randomly chose 45 triplets with responses from between the two models. The final dataset contains 132 question-answer-context triplets, along with the human ratings for CF, CR and RA from independent labellers. We name this dataset UniqueQAC and make it available for research purposes.
E.6 Creating a dataset for clinical harm, helpfulness and inappropriateness
To support Claim (3), we start with UniqueQAC and ask two ophthalmic surgeons to label the answers for the following axes:
  1. 1.
    Clinical Harm: Is the response harmful?
  2. 2.
    Helpfulness: Is the response helpful?
  3. 3.
    Appropriateness: Is there inappropriate or incorrect content?
After the surgeons resolved disagreements, we combined UniqueQAC with the clinician ratings for harmfulness, helpfulness, and inappropriateness. This resulted in a dataset where most responses exhibited no harm.
To balance the dataset for each of the three categories, we replaced responses from the clinical QA system with those from a clinician who provided potentially harmful, unhelpful, and inappropriate responses to the patient questions. The final dataset, containing 132 question-answer-context triplets, is named ClinicalQAC (pun intended) and is released for research purposes. Figure 16 illustrates the dataset proportions.
Figure 16: ClinicalQAC: Proportions of different categories in the harmfulness, helpfulness and inappropriateness axes.
Appendix F Experimental details
We provide information on training and hyperparameter tuning details in this section.
Random Forest Classifier
We implement a random forest classifier using Scikit-learn. We perform grid on the parameters
SVM
We implement an SVM using Scikit-learn.
Gaussian Naive Bayes
We implement an Gaussian Naive Bayes using Scikit-learn.
Neural Network
We implement a simple neural network using PyTorch.
Appendix G Inter-annotator agreements
The initial set of clinical assessments included five axes.
  1. 1.
    Inappropriateness: Is there inappropriate or incorrect content?
  2. 2.
    Intent: Does it address the question intent?
  3. 3.
    Helpfulness: How helpful is the answer to the user?
  4. 4.
    Extent of Harm: In this clinical context, what is the extent of possible harm?
  5. 5.
    Likelihood of Harm: In this clinical context, what is the likelihood of possible harm?
We observed that "Intent" and "Helpfulness" were quite interdependent and so we combined them into the broad category of Helpfulness. We observed similar interdependence between Extent and Likelihood of harm and thus combined them into Harmfulness.
Metric Value
Is there inappropriate or incorrect content? 0.65
Does it address the intent of the question? 0.93
How helpful is the answer to the user? 0.77
In this clinical context, what is the extent of possible harm? 0.90
In this clinical context, what is the likelihood of possible harm? 0.95
Table 4: Inter-annotator agreement on clinical axes
Appendix H Additional LLM Performance Analysis on ASTRID Metrics
To provide a more detailed view of model performance across the ASTRID metrics, Figures 17, 18, and 19 present Accuracy, F1-score, Precision, and Recall for each model across Conversational Faithfulness, Context Relevance, and Refusal Accuracy.
Figure 17: Model-wise performance on Conversational Faithfulness across Accuracy, F1-score, Precision, and Recall.
Figure 18: Model-wise performance on Context Relevance across Accuracy, F1-score, Precision, and Recall.
Figure 19: Model-wise performance on Refusal Accuracy across Accuracy, F1-score, Precision, and Recall.
Model Provider Date Temperature Top p Token Limit
llama3-3-70b-instruct AWS 12/02/2025 0.1 0.9 200
llama3-8b-instruct AWS 12/02/2025 0.1 0.9 200
mistral-7b-instruct AWS 12/02/2025 0.1 0.9 200
mistral-large AWS 12/02/2025 0.1 0.9 200
gemini-2.0-flash Google 12/02/2025 0.1 0.9 200
claude-3.5-sonnet Google 12/02/2025 0.1 0.9 200
gpt-4o Azure 12/02/2025 0.1 0.9 200
gpt-o3-mini Azure 12/02/2025 N/A N/A 1000
Table 5: Inference details for all evaluated LLMs.